OpenStudy (anonymous):

please solve (d^2+1)y=e^-1+cosx+x^3+e^xsinx

OpenStudy (zzr0ck3r):

for what?

ganeshie8 (ganeshie8):

\[\large (D^2+1)y=e^{-1+\cos x}+x^3+e^x\sin x\] like this ?

OpenStudy (zzr0ck3r):

lol good luck with that

OpenStudy (anonymous):

(D^2 +1)y=e^−x+cosx + x^3 +e^x sinx

ganeshie8 (ganeshie8):

\[\large (D^2+1)y=e^{-x}+\cos x+x^3+e^x\sin x\] like this ?

OpenStudy (anonymous):

yes my friend u r right. pls help me with ths

OpenStudy (anonymous):

Hi ganesh r u thr?

ganeshie8 (ganeshie8):

I think general solution should be straight forward right ? for particular solution you may use superposition

ganeshie8 (ganeshie8):

\(r^2 + 1 = 0 \implies r = \pm i \) the general solution would be : \(y = c_1 \cos x + c_2 \sin x\) use superposition for finding the particular solution

ganeshie8 (ganeshie8):

for example : \(e^{-x} \) gives you \(\large \dfrac{e^{-x}}{(-1)^2+1} = \dfrac{e^{-x}}{2}\)

ganeshie8 (ganeshie8):

similarly find the particular solutions for each of the other terms : cosx, x^3 and e^xsinx then add them up

OpenStudy (zzr0ck3r):

yep im lost